OpenStudy (anonymous):

prove sin 3 theta = 3 sin theta - 4 sin ^3 theta

OpenStudy (anonymous):

sin(3t) = sin(2t + t) = sin(2t)cos(t) + cos(2t)sin(t) helps ?